OpenStudy (clamin):

ONE MEDAL!! Determine the area of the shaded region.The radius of the larger circle is 6m, and the radius of the smaller circle is 4m.

OpenStudy (anonymous):

\[\pi 6^{2}- \pi 4^{2} = 20 \pi \]